tan-1( 2x/1-x^2) + cot-1 1-x^2/2x^2) = pie/3


plz proove

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

tan-1(2x/(1-x2)=2tan-1x

cot-1((1-x2)/2x2)=tan-1(2x2/(1-x2))

use the formula: tan-1a + tan-1b =tan-1((a+b)/(1-ab))

ankit dewangan
12 years ago

now i have only 20 days to my board exams.... suggest some tricks and time management for phy. and chem..

regards